solve system if possible
27x-3y=36
9x-12=y

Answer by checkley77(12844)

27x-3y=36
9x-12=y MULTIPLY BY 3
27X-3Y=36
THESE 2 EQUATIONS ARE THE SAME.
THEREFORE THERE IS NO UNIQUE SOLUTION FOR X & Y.